Home‎ > ‎Reports‎ > ‎Crew097‎ > ‎

day11


Crew 97 Log Book for January 11, 2011

Photos of the day





Commander's Check-In Report

Judah Epstein Reporting

Crew Physical Status:

Very Good. Increased crew morale.


Time Departed/Returned from EVA:

EVA #20: 09:54 / 10:27. Jim.

EVA #21: 12:46 / 13:40. Nathan, Jim.

EVA #22: 15:08 / 16:25. Jim, Nathan.


Brief Narrative of Field Mission Results:

EVAs today consisted of Rock Varnish collection, geocache correction, and Satellite data comparison collection.


EVA Data/Interpretations:

EVA #20: Location of geocache was verified with correct coordinates.

EVA #21: Rock Varnish samples were successfully collected based upon returning to previously determined location.

EVA #22: Satellite data comparison photos were taken from the top of Olympus Mons.


Engineering/Hab Maintenance:

Hot water heater was repaired with DG bringing new electric control panel and re-wiring connections.

EVA Pack #3 was repaired and all 6 packs are functional.

Webcams repaired and all webcams now operational.


Report Transmission Schedule:

CMDR Check-in, ENGR, CMDR Report, Photos, 3 EVA Reports, 3 Science Reports


Plans for Tomorrow:

- Cooking/Cleaning - Jim/Lucinda

- 07:30 - Breakfast

- 8:00 - Exercise

- 8:20-9:30 - Psychophysiological crew analysis

- Engineering checks

- EVA #23: Mobility Study with everyone

- 11:30 Lunch

- 12:00-1:10 - Psychophysiological crew analysis

- EVA #24: Water Gully Analysis

- EVA #25: Satellite Data Comparison

- EVA #26: Rover testing

- 17:30 - Dinner

- Psychophysiological crew analysis

- Reports, Food Study

- Work on experiments

- Showers - Amanda/Jim


Inventory:

Kitchen supplies inventoried.


Support Requested:

Delivery of dehydrated food.

Recommendation for additional defog spray. Polo Antibeschlag-Spray from Fairchild Sports Switzerland GmbH seems to work best. It is recommended that MDRS provide more spray or for crews to bring their own. The current container at MDRS is nearly depleted.


Miscellaneous:

Kitchen extensively cleaned and organized.


Commander's Report

Judah Epstein Reporting

Another exciting day on Mars was successfully completed. The crew's morale has greatly increased with dividing up of snacks and therefore snacks more readily available, repair of hot water heater so that now crew-members will take showers, and Tonya making a large batch of chocolate chip cookies.

After breakfast and exercise the crew began their various tasks and experiments for the day. The Crew Engineer worked on engineering responsibilities and repairs as well as participating on EVAs including collecting samples for his Rock Varnish Study to test for signs of life. The Automated Exploration Scientist was busy with various tasks such as troubleshooting communication errors with the Sandstorm Rover. The Crew Geologist organized, cleaned, and inventoried the kitchen as well as preparations for tomorrow's Water Gully EVA. The Usability Engineer focused on psychophysiological experiments and testing. The Crew Commander did extensive testing on water samples from the slow sand filters including bacteria testing, and also worked on various engineering repairs. And as usual, the Executive Officer was the lead for EVA exploration.

Tomorrow will consist of another full day of experiments such as extended psychophysiological analysis and a full day of four planned EVAs including the EVA Mobility Study with full crew-member participation.

Science Report 1

Judah Epstein Reporting

The purpose of this project is to explore the effectiveness of a slow sand sedimentary water filtration system in a planetary environment. To simulate a planetary environment the experiment is conducted at the Mars Desert Research Station, a self-sustaining planetary research experiment and simulation expedition.

Background information regarding the project abstract, purpose, goals, and specific experiment details and information are included in the Science Report from January 6.

Data collected from Jan 10 and prior is available in this spreadsheet.



Updates on results:

Filter 1 (Martian Simulant) and Filter 2 (Commercial Sand) are not achieving optimal results as the filter seems clogged with extremely low hydraulic conductivity. Therefore not enough effluent water was produced to take measurements with the Qanta probe (Temp, SpC, DO, pH, TDS, DO%, ORP).

Turbidity results: Turbidity of effluent from all three slow sand filters has generally been increasing. Turbidity of effluent from the wetland, greywater, and HAB potable is remaining generally consistent.


Science Report 2

Tonya Thompson Reporting

I performed 3 psychophysiological reading sessions on each crewmember today. Each session lasted 3 minutes.

The crewmember entered my stateroom, sat in a chair, placed finger sensors on their 3 middle fingers, and watched a white image projected onto a wall for 3 minutes. During that time, I recorded their heart rate variability (hrv) and skin conductance level (scl). The data for each session was saved to the hard drive of a desktop pc in my stateroom.

I also had each crewmember fill out a Cohen’s Perceived Stress Scale questionnaire at the end of the evening after dinner.

Science Report 2

Tonya Thompson Reporting

I performed 3 psychophysiological reading sessions on each crewmember today. Each session lasted 3 minutes.

The crewmember entered my stateroom, sat in a chair, placed finger sensors on their 3 middle fingers, and watched a white image projected onto a wall for 3 minutes. During that time, I recorded their heart rate variability (hrv) and skin conductance level (scl). The data for each session was saved to the hard drive of a desktop pc in my stateroom.

I also had each crewmember fill out a Cohen’s Perceived Stress Scale questionnaire at the end of the evening after dinner.


Science Report 3

Lucinda Land Reporting

Continuing issues with connection between Sandstorm and the router (Maxnet) were ongoing. Yesterday we had new problems with connecting outside. It was an odd occurrence since we had connected to her outside before under the same conditions. We then brought her inside to check connection to Maxnet and had no success. This was even more odd because we had never had any issues with connecting her while inside the Hab.

The problem, it seems, was not with Sandstorm, or Maxnet. Both were functioning as normal. The crew commander could connect to Maxnet on his computer with out issues. Something had changed with the configuration on my computer that was not allowing Maxnet to connect properly.

I have sent emails to the remote team on standby at NASA Ames and part of our NorCal Mars Society chapter. They are troubleshooting the issue this evening and we hope to have her up and running by tomorrow to finish the distance and route range tests.

In the meantime, I will start to use the Gigapan camera tonight and tomorrow morning with 360 degree image of the inside of the Hab. With Sandstorm working later again, we'll take her out for a final spin and use photo capture and take further Gigapan imagery of the local Geology.


Engineering Report

Nathan Wong Reporting

Generator/Electricity:

Kitty(Diesel Generator) (Used? Yes/No): Yes

Honey (Gas Generator) (Used? Yes/No): No

Battery Charge Level:

Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off): Power worked well today


Fuel Status:

Diesel (Full/0.75/0.5/0.25/Empty): 0.50

Propane Tank (Full/0.75/0.5/0.25/Empty): 0.51

Gasoline Tank (Full/0.75/0.5/0.25/Empty): 3 Gas cans full

Notes/Comments:


Oil:

Kitty Oil Quantity (# of Quarts in storage): 3

Honey Oil Quantity (# of Quarts in storage): 2

ATV Oil Quantity (# of Quarts in storage): 2

Car Oil Quantity (# of Quarts in storage): none


Water Status:

Outside Potable Water Tank Level (inches from bottom): 40%

Trailer Potable Water Tank Level (inches from bottom): 75%

Hab Potable Water Tank Level (inches from bottom): empty

Potable Water Meter Reading: 40770.0

In to GreenHab Meter Reading:

Notes/Comments:

I will try to heat the new water line tomorrow and see if that fixes the problem.

The toilet line is slow to flush. DG thinks the line is partially frozen, I will check the heater tape tomorrow.


GreenHab:

Condition of plants in Tank 1: None

Condition of plants in Tank 2: 3 plants

Condition of plants in Tank 3: 7 plants

Greenhab Notes/Comments: All plants are good.


Observatory:

Telescope Used? (Yes/No): No

Observatory Notes/Comments: Observatory not used


Transportation:

Hab Car (NOT RENTAL)
Used (yes/no): No
Oil Checked (yes/no): No
Tire Status: good
Odometer at end of day: 123085
Notes/Comments on Hab Car:

ATV 1 (Opportunity):
Used: Yes
Oil Checked: No
Fuel Consumed: Yes
Tire Status: Inflated
Comments:

ATV 2 (Spirit):
Used: Yes
Oil Checked: No
Fuel Consumed: Yes
Tire Status: Inflated
Comments:

ATV 3 (Viking I):
Used: No
Oil Checked: No
Fuel Consumed: No
Tire Status: Inflated
Comments:

ATV 4 (Viking II):
Used: Yes
Oil Checked: No
Fuel Consumed: Yes
Tire Status: Inflated
Comments:


Heating and Ventilation:
Thermostat setting upstairs (Farenheit): off
Actual temperature upstairs (Farenheit): 72

Thermostat setting downstairs (Farenheit): 55
Actual temperature downstairs (Farenheit): 57


Computers/Networking Infrastructure:

Notes/comments on internet/computers?: All webcams are now functioning properly.


General Engineering Notes/Comments:

-Pack three is now repaired. All packs are operational.

- Two packs need new fuse holders. Right now the fuse is being bypassed in order to make the pack operational.

-Hot water heater is now working

EVA 20 Report

Jim Crowell Reporting

Time: 09:54 / 10:27

Crew Members: Jim Crowell

Site Location: West of Hab

Transit Mode: Walking

Objectives: To find the real location of the MDRS Geocache

Results: Location found, and real GPS coordinates recorded.

Lessons Learned: When recording GPS coordinates by hand, ensure that the numbers are correct.

EVA 21 Report

Nathan Wong Reporting

Time:

Enter Airlock: 12:46
Exit Airlock: 1:40

Crew Members: Nathan Wong and Jim Crowell

Site Location: East of the Hab

Transit Mode: ATV

Objectives: Collect rock varnish samples.

Results: Two rock varnish samples and one soil sample were collected and cataloged. These samples will be taken back and analaysis will be performed to test for signs of life.

Lessons Learned: Collecting samples while in full suit was not easy, especially opening up the sterilized bags.

EVA 22 Report

Jim Crowell Reporting

Time: 15:08 / 16:25

Crew Members: Jim Crowell, Nathan Wong

Site Location: Olympus Mons

Transit Mode: ATVs / Walking

Objectives: To collect panoramic images from the top of Olympus Mons

Results: Rode ATVs out to the mountain, walked to the top, and took five panoramas.

Lessons Learned: On hiking EVAs (even if it's chilly out), be sure to wear light clothes. The extra effort will make you warmer than originally anticipated.

Cook's Report

Amanda Damptz Reporting

1a) Was today a cooking day? Yes

1b) Was today a special day celebrated at the main meal with special food or activity? No special day but Tonya made Oatmeal Chocolate Chip Cookies.

2) List the foods served at today's main meal giving full names of each.

Pancakes (Blueberry pancake boxed mix and recipe),
Bacon (shelf stable bacon)
Whole egg mix
Potatoes (made into hash browns with pepper and basil)
Bear Creek Vegetable Beef Soup
Oatmeal chocolate chip cookies

3) List any main meal foods not finished by the crew (leftovers stored for later use or discarded) Vegetable Beef Soup

4) Recipes for anything more complicated than following package directions exactly, or rehydrating.
    Pancake Recipe
1 1/2 cups all-purpose flour
3 1/2 teaspoons baking powder
1 teaspoon salt
1 tablespoon white sugar
1 1/4 cups milk (reconstituted)
1 egg (reconstituted)
3 tablespoons shelf stable butter, melted

Whisk all ingredients together, let stand a few minutes to thicken, bake pancakes on greased frying pan or griddle.

Cookies
No recipe

Hash brown potatoes
No recipe

5) Comments/questions for the food study investigators: NONE